# Sup-norm bounds for Jacobi cusp forms

### v2

@@ Abstract.
In this article, we give bounds for the natural invariant norm of cusp forms of real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. Using the representation of Jacobi cusp forms of integral weight $k$ and index $m$ for the modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$ as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions, we derive bounds for the natural invariant norm of these Jacobi cusp forms. More specifically, letting $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the complex vector space of Jacobi cusp forms under consideration and $\|\cdot\|_{\mathrm{Pet}}$ the pointwise Petersson norm on $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we prove that for given $\epsilon>0$, the bound
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}m^{1+\epsilon}\big{)}$
holds for any $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, where the implied constant depends only on the choice of $\epsilon>0$.
@@ 1.1. Background
In general, bounds for automorphic forms and for their Fourier coefficients represent an area of great interest in number theory. More specifically, we mention in this respect the results of [FJK16], where J. Friedman, J. Jorgenson, and J. Kramer obtained optimal sup-norm bounds on average for cusp forms of even weight $k$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. These bounds turn out to be uniform with respect to the subgroup $\Gamma$. Moreover, in [FJK19], effective versions for these sup-norm bounds are given. With regard to sup-norm bounds for individual Hecke eigenforms of large level, we mention, for example, the results by V. Blomer and R. Holowinsky in [BH11].
@@
So far, less attention has been devoted to the study of sup-norm bounds for Jacobi cusp forms. The first comprehensive study of Jacobi forms was undertaken by M. Eichler and D. Zagier in [EZ85]. Subsequently, various authors have built on their work. In contrast to their analytical approach, a geometrical approach to the theory of Jacobi forms was given by J. Kramer in [Kr91].
@@
Let $k,m$ be positive integers. A Jacobi form of weight $k$ and index $m$ for the modular group $\Gamma_{0}\coloneqq\mathrm{SL}_{2}(\mathbb{Z})$ is a holomorphic function on the product $\mathbb{H}\times\mathbb{C}$ of the upper half-plane $\mathbb{H}$ with the complex plane $\mathbb{C}$ having a suitable transformation behaviour with respect to $\Gamma_{0}$ and vanishing “at infinity”. We denote the complex vector space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$ by $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. The pointwise Petersson norm of a Jacobi form $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is then defined by
$\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\coloneqq|f(\tau,z)|^{2}\,\mathrm{Im}(\tau)^{k}\,e^{-\frac{4\pi m\mathrm{Im}(z)^{2}}{\mathrm{Im}(\tau)}}\qquad(\tau\in\mathbb{H},\,z\in\mathbb{C}).$
@@
Let $F$ be a Siegel cusp form of weight $k$ for the Siegel modular group $\mathrm{Sp}_{4}(\mathbb{Z})$, and let $\{f_{m}\}_{m\geq 1}$ be the set of Jacobi forms appearing in the Fourier–Jacobi expansion of $F$, i. e., $f_{m}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. Then, for any $\epsilon>0$, W. Kohnen proved the following sup-norm bound for the pointwise Petersson norm of $f_{m}$ in [Ko93]
(1) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k}{2}-\frac{2}{9}+\epsilon}\big{)},$
@@
where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$. Motivated by the Ramanujan–Petersson conjecture, W. Kohnen then conjectured the bound
$\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k-1}{2}+\epsilon}\big{)},$
@@
where the implied constant depends on the Siegel cusp form $F$ and on the choice of $\epsilon>0$.
@@
More recently, P. Anamby and S. Das established in [AD23] a general sup-norm bound for the pointwise Petersson norm of $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, i. e., for which we have
@@
$\int\limits\limits_{\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|^{2}_{\mathrm{Pet}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}=1\qquad(\tau=\xi+i\eta,\,z=x+iy).$
Their bound is (see Theorem 1.4 in[AD23])
(2) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k\,m\big{)},$
@@ 1.2. Main results
The goal of this article is to provide new sup-norm bounds for the pointwise Petersson norm for Jacobi forms of integral weight $k$ and integral index $m$ for $\Gamma_{0}$, which are normalized with respect to the Petersson inner product. The main result in this respect is given in Theorem 4.4 and states for $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is normalized with respect to the Petersson inner product, that
(3) $\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|f(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}k^{\frac{3}{4}}\,m^{1+\epsilon}\big{)},$
where the implied constant depends only on the choice of $\epsilon>0$. For the proof, we essentially use the representation of the Jacobi cusp forms under consideration as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions; we then need to derive bounds for the pointwise Petersson norms of these functions to arrive at our result. Comparing our bound (3) with the bound (2) by P. Anamby and S. Das, we realize an improvement with regard to the polynomial growth in $k$ by a factor of $k^{\frac{1}{4}}$, while the polynomial growth of the bound (2) in $m$ is better by a factor of $m^{\epsilon}$ (for any $\epsilon>0$).
In order to be able to derive our bound (3), we need sup-norm bounds for the pointwise Petersson norm of cusp forms of positive real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{R})$. Such bounds have been derived for finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by Steiner in [St16]. Although these bounds are uniform with respect to the index of the finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$, there are mild restrictions for a direct application of the bounds derived in [St16] to our setting. On the other hand, such bounds could also be derived from [FJK16] with some extra work. However, we provide here new, alternative proofs for the results of [FJK16] applying to any positive real weight $k$ and any character $\chi$ by using the Bergman kernel for the modular curve associated to $\Gamma$.
@@
As a second main result of this article, we establish in Theorem 3.3, assuming that $k\in\mathbb{R}_{\geq 5}$, for $\Gamma$ being cocompact without elliptic elements the bound
$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k);$
moreover, for $\Gamma$ being cofinite, we give the bound
$\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)},$
where the implied constants depend only on the Fuchsian subgroup $\Gamma$. Due to the relation (4), these results reprove the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Based on these results, we are then also able to prove the uniformity of the above bounds with respect to the subgroup $\Gamma$ in Theorem 3.5.
@@ 1.3. Outline
Let us briefly outline the contents of this article. In the subsequent, second section we collect all the necessary prerequisites for the sequel of the paper. In particular, we introduce the definitions of cusp forms and Jacobi cusp forms together with their respective (pointwise) Petersson inner products. Furthermore, we define the Bergman kernel for modular curves and state its basic properties. We close the section by recalling asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles on compact complex Kähler manifolds due to [Bo96], which are crucial in the derivation of the bound (3).
The third section is devoted to the revisiting of the sup-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Here, the proofs of Theorem 3.3 and Theorem 3.5 are provided.
In the fourth section, the bound (3) is proven in Theorem 4.4. In addition to some straightforward bounds established for the relative $\mathrm{L}^{2}$-norm of classical Jacobi theta functions, the above mentioned asymptotics of Bergman kernels associated to tensor powers of holomorphic line bundles corresponding to these Jacobi theta functions are crucial in the derivation of the proof of Theorem 4.4.
@@ 2.2. Quotient space
which is compact; we note that $\mathcal{F}_{Y}=\mathcal{F}_{\Gamma}$, if $\Gamma$ is cocompact. We choose $0<m_{Y}<M_{Y}$ such that for all $\tau\in\mathcal{F}_{Y}$ the inequalities
@@
(8) $r_{\Gamma}\coloneqq\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{\Gamma},\,\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\bigg{\}}.$
@@ 2.4. Counting function
$\mathcal{N}_{\Gamma}(\tau;\rho)\coloneqq\bigg{\{}\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\bigg{)}\,\bigg{|}\,\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\leq\rho\bigg{\}}.$
@@ 2.5. Jacobi forms
For $k,m\in\mathbb{N}$, we let $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, i. e., the space of holomorphic functions $f\colon\mathbb{H}\times\mathbb{C}\rightarrow\mathbb{C}$, which have the transformation behaviour
$f\Big{(}\frac{a\tau+b}{c\tau+d},\frac{z+\lambda\tau+\mu}{c\tau+d}\Big{)}(c\tau+d)^{-k}\exp\Big{(}2\pi im\Big{(}\lambda^{2}\tau+2\lambda z-\frac{c(z+\lambda\tau+\mu)^{2}}{c\tau+d}\Big{)}\Big{)}=f(\tau,z)$
@@
$f(\tau,z)=\sum\limits_{\begin{subarray}{c}n\in\mathbb{N},\,r\in\mathbb{Z}\\ 4mn-r^{2}>0\end{subarray}}c(n,r)\,q^{n}\zeta^{r}\qquad(q=e^{2\pi i\tau},\,\zeta=e^{2\pi iz}).$
Given $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we define
$\|f(\tau,z)\|_{\mathrm{Pet}}^{2}\coloneqq|f(\tau,z)|^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\qquad(\tau=\xi+i\eta,\,z=x+iy),$
which defines a $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariant function on $\mathbb{H}\times\mathbb{C}$ called the pointwise Petersson norm of $f$.
@@
(12) $\langle f_{1},f_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits\limits_{\mathcal{D}_{\Gamma_{0}}}f_{1}(\tau,z)\overline{f_{2}(\tau,z)}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}\qquad(f_{1},f_{2}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})).$
For $f\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, one has the decomposition
(13) $f(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the function $\varphi_{\mu}$ is a cusp form of weight $(k-\frac{1}{2})$ for the finite index subgroup $\Gamma_{1}\coloneqq\Gamma_{0}(4m)$ of $\Gamma_{0}$, and $\vartheta_{\mu,m}$ is the Jacobi theta function
@@
As we will see below, the theta functions $\vartheta_{\mu,m}$ ($\mu=0,\ldots,2m-1$) arise for fixed $\tau\in\mathbb{H}$ as global sections of a suitable line bundle on the elliptic curve associated to $\tau$. In fact, it is shown in Theorem 5.1 of [EZ85] that the decomposition (14) gives rise to the isomorphism
@@
$f_{1}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,1}(\tau)\vartheta_{\mu,m}(\tau,z)\quad\text{and}\quad f_{2}(\tau,z)=\sum_{\mu=0}^{2m-1}\varphi_{\mu,2}(\tau)\vartheta_{\mu,m}(\tau,z)$

### v3

@@ Abstract.
In this article, we give $L^{\infty}$-norm bounds for the natural invariant norm of cusp forms of real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. Using the representation of Jacobi cusp forms of integral weight $k$ and index $m$ for the modular group $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$ as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions, we derive $L^{\infty}$-norm bounds for the natural invariant norm of these Jacobi cusp forms. More specifically, letting $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the complex vector space of Jacobi cusp forms under consideration and $\|\cdot\|_{\mathrm{Pet}}$ the pointwise Petersson norm on $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we prove that for $k\in\mathbb{Z}_{\geq 5}$ and $m\in\mathbb{Z}_{\geq 1}$, and a given $\epsilon>0$, the $L^{\infty}$-norm bound
$\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k\,m^{\frac{7}{4}+\epsilon}\big{)}$
holds for any $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is $L^{2}$-normalized with respect to the Petersson inner product, where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$.
@@ 1.1. Background
In general, bounds for automorphic forms and for their Fourier coefficients represent an area of great interest in number theory. More specifically, we mention in this respect the results of [FJK16], where J. Friedman, J. Jorgenson, and J. Kramer obtained optimal $L^{\infty}$-norm bounds on average for cusp forms of even weight $k$ for any cofinite Fuchsian subgroup $\Gamma\subset\mathrm{SL}_{2}(\mathbb{R})$. These $L^{\infty}$-norm bounds turn out to be uniform with respect to the subgroup $\Gamma$. Moreover, in [FJK19], effective versions for these $L^{\infty}$-norm bounds are given. With regard to $L^{\infty}$-norm bounds for individual Hecke eigenforms of large level, we mention, for example, the results by V. Blomer and R. Holowinsky in [BH11].
@@
So far, less attention has been devoted to the study of $L^{\infty}$-norm bounds for Jacobi cusp forms. The first comprehensive study of Jacobi forms was undertaken by M. Eichler and D. Zagier in [EZ85]. Subsequently, various authors have built on their work. In contrast to their analytical approach, a geometrical approach to the theory of Jacobi forms was given by J. Kramer in [Kr91].
@@
Let $k,m$ be positive integers. A Jacobi form of weight $k$ and index $m$ for the modular group $\Gamma_{0}\coloneqq\mathrm{SL}_{2}(\mathbb{Z})$ is a holomorphic function on the product $\mathbb{H}\times\mathbb{C}$ of the upper half-plane $\mathbb{H}$ with the complex plane $\mathbb{C}$ having a suitable transformation behaviour with respect to $\Gamma_{0}$ and vanishing “at infinity”. We denote the complex vector space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}$ by $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. The pointwise Petersson norm of a Jacobi form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ is then defined by
$\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}\coloneqq|\phi(\tau,z)|^{2}\,\mathrm{Im}(\tau)^{k}\,e^{-\frac{4\pi m\mathrm{Im}(z)^{2}}{\mathrm{Im}(\tau)}}\qquad(\tau\in\mathbb{H},\,z\in\mathbb{C}).$
@@
Let $F$ be a Siegel cusp form of weight $k$ for the Siegel modular group $\mathrm{Sp}_{4}(\mathbb{Z})$, and let $\{\phi_{m}\}_{m\geq 1}$ be the set of Jacobi cusp forms appearing in the Fourier–Jacobi expansion of $F$, i. e., $\phi_{m}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$. Then, for any $\epsilon>0$, W. Kohnen proved the following $L^{\infty}$-norm bound for the pointwise Petersson norm of $\phi_{m}$ in [Ko93]
(1) $\|\phi_{m}\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k}{2}-\frac{2}{9}+\epsilon}\big{)},$
@@
where the implied constant depends on the Siegel cusp form $F$ and the choice of $\epsilon>0$. Motivated by the Ramanujan–Petersson conjecture, W. Kohnen then conjectured the $L^{\infty}$-norm bound
$\|\phi_{m}\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi_{m}(\tau,z)\|_{\mathrm{Pet}}=O_{F,\epsilon}\big{(}m^{\frac{k-1}{2}+\epsilon}\big{)},$
@@
where the implied constant depends on the Siegel cusp form $F$ and the choice of $\epsilon>0$.
@@
More recently, P. Anamby and S. Das established in [AD23] a general $L^{\infty}$-norm bound for the pointwise Petersson norm of any Jacobi cusp form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, which is $L^{2}$-normalized with respect to the Petersson inner product, i. e., for which we have
@@
$\|\phi\|^{2}_{L^{2}}=\int\limits\limits_{\Gamma_{0}\ltimes\mathbb{Z}^{2}\backslash\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|^{2}_{\mathrm{Pet}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}=1\qquad(\tau=\xi+i\eta,\,z=x+iy).$
Their $L^{\infty}$-norm bound is (see Theorem 1.4 in[AD23])
(2) $\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\epsilon}\big{(}(k\,m)^{1+\epsilon}\big{)},$
@@ 1.2. Main results
The goal of this article is to provide $L^{\infty}$-norm bounds for the pointwise Petersson norm for Jacobi forms of weight $k\in\mathbb{Z}_{\geq 5}$ and index $m\in\mathbb{Z}_{\geq 1}$ for $\Gamma_{0}$, which are $L^{2}$-normalized. The main result in this respect is given in Theorem 4.4 and states for any $L^{2}$-normalized Jacobi cusp form $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ the $L^{\infty}$-norm bound
(3) $\|\phi\|_{L^{\infty}}=\sup_{(\tau,z)\in\mathbb{H}\times\mathbb{C}}\|\phi(\tau,z)\|_{\mathrm{Pet}}=O_{\Gamma_{0},\epsilon}\big{(}k\,m^{\frac{7}{4}+\epsilon}\big{)}$
holds, where the implied constant depends on $\Gamma_{0}$ and the choice of $\epsilon>0$. For the proof, we essentially use the representation of the Jacobi cusp forms under consideration as linear combinations of modular forms of weight $k-\frac{1}{2}$ for some congruence subgroup of $\Gamma_{0}$ (depending on $m$) and suitable Jacobi theta functions; we then need to derive bounds for the $L^{\infty}$-norms of these functions to arrive at our result. Comparing our $L^{\infty}$-norm bound (3) with the $L^{\infty}$-norm bound (2) by P. Anamby and S. Das, shows that the polynomial growth in $k$ is the same, while the polynomial growth of their bound in $m$ is better by a factor of $m^{\frac{3}{4}}$.
In order to be able to derive our $L^{\infty}$-norm bound (3), we need $L^{\infty}$-norm bounds for the pointwise Petersson norm of cusp forms of positive real weight $k$ and character $\chi$ for any cofinite Fuchsian subgroup $\Gamma$ of $\mathrm{SL}_{2}(\mathbb{R})$. Such $L^{\infty}$-norm bounds have been derived for finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$ by Steiner in [St16]. Although these $L^{\infty}$-norm bounds are uniform with respect to the index of the finite index subgroups of $\mathrm{SL}_{2}(\mathbb{Z})$, there are mild restrictions for a direct application of the $L^{\infty}$-norm bounds derived in [St16] to our setting. On the other hand, such $L^{\infty}$-norm bounds could also be derived from [FJK16] with some extra work. However, we provide here new, alternative proofs for the results of [FJK16] applying to any positive real weight $k$ and any character $\chi$ by using the Bergman kernel for the modular curve associated to $\Gamma$.
@@
As a second main result of this article, we establish in Theorem 3.3, assuming that $k\in\mathbb{R}_{\geq 5}$, for $\Gamma$ being cocompact without elliptic elements the $L^{\infty}$-norm bound
$\|B_{k,\chi}\|_{L^{\infty}}=\sup_{\tau\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}(k);$
moreover, for $\Gamma$ being cofinite, we give the $L^{\infty}$-norm bound
$\|B_{k,\chi}\|_{L^{\infty}}=\sup_{z\in\mathbb{H}}\|B_{k,\chi}(\tau,\tau)\|_{\mathrm{Pet}}=O_{\Gamma}\big{(}k^{\frac{3}{2}}\big{)},$
where the implied constants depend only on the Fuchsian subgroup $\Gamma$. Due to the relation (4), these results reprove the ${L^{\infty}}$-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Based on these results, we are then also able to prove the uniformity of the above ${L^{\infty}}$-norm bounds with respect to the subgroup $\Gamma$ in Theorem 3.5.
@@ 1.3. Outline
Let us briefly outline the contents of this article. In the subsequent, second section we collect all the necessary prerequisites for the sequel of the paper. In particular, we introduce the definitions of cusp forms and Jacobi cusp forms together with their respective pointwise Petersson norms. Furthermore, we define the Bergman kernel for modular curves and state its basic properties.
The third section is devoted to the revisiting of the ${L^{\infty}}$-norm bounds on average obtained in [FJK16], but now for any real weight $k\in\mathbb{R}_{\geq 5}$ and any character $\chi$. Here, the proofs of Theorem 3.3 and Theorem 3.5 are provided.
In the fourth section, the ${L^{\infty}}$-norm bound (3) is proven in Theorem 4.4. In order to establish this result, we combine the results of the third section applied to modular forms of half-integral weight with $L^{\infty}$-norm bounds of classical Jacobi theta functions, which are derived by directly estimating their defining series expansions.
@@ 2.2. Quotient space
which is compact. For $Y$ being sufficiently large, we note that $\mathcal{F}_{Y}=\mathcal{F}_{\Gamma}$, if $\Gamma$ is cocompact. We choose $0<m_{Y}<M_{Y}$ such that for all $\tau\in\mathcal{F}_{Y}$ the inequalities
@@
(8) $r_{\Gamma}\coloneqq\inf\bigg{\{}\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\,\bigg{|}\,\tau\in\mathcal{F}_{\Gamma},\,\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\cup\{\mathrm{id}\}\bigg{)}\bigg{\}}.$
@@ 2.4. Counting function
$\mathcal{N}_{\Gamma}(\tau;\rho)\coloneqq\bigg{\{}\gamma\in\Gamma\setminus\bigg{(}\bigcup_{j=1}^{s}\Gamma_{p_{j}}\cup\bigcup_{j=1}^{t}\Gamma_{e_{j}}\cup\{\mathrm{id}\}\bigg{)}\,\bigg{|}\,\mathrm{dist}_{\mathrm{hyp}}(\tau,\gamma\tau)\leq\rho\bigg{\}}.$
@@ 2.5. Jacobi forms
For $k,m\in\mathbb{N}$, we let $J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$ denote the space of Jacobi cusp forms of weight $k$ and index $m$ for $\Gamma_{0}=\mathrm{SL}_{2}(\mathbb{Z})$, i. e., the space of holomorphic functions $\phi\colon\mathbb{H}\times\mathbb{C}\rightarrow\mathbb{C}$, which have the transformation behaviour
$\phi\Big{(}\frac{a\tau+b}{c\tau+d},\frac{z+\lambda\tau+\mu}{c\tau+d}\Big{)}(c\tau+d)^{-k}\exp\Big{(}2\pi im\Big{(}\lambda^{2}\tau+2\lambda z-\frac{c(z+\lambda\tau+\mu)^{2}}{c\tau+d}\Big{)}\Big{)}=\phi(\tau,z)$
@@
$\phi(\tau,z)=\sum\limits_{\begin{subarray}{c}n\in\mathbb{N},\,r\in\mathbb{Z}\\ 4mn-r^{2}>0\end{subarray}}c(n,r)\,q^{n}\zeta^{r}\qquad(q=e^{2\pi i\tau},\,\zeta=e^{2\pi iz}).$
Given $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, we define
$\|\phi(\tau,z)\|_{\mathrm{Pet}}^{2}\coloneqq|\phi(\tau,z)|^{2}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\qquad(\tau=\xi+i\eta,\,z=x+iy),$
which defines a $\Gamma_{0}\ltimes\mathbb{Z}^{2}$-invariant function on $\mathbb{H}\times\mathbb{C}$ called the pointwise Petersson norm of $\phi$.
@@
(12) $\langle\phi_{1},\phi_{2}\rangle_{\mathrm{Pet}}\coloneqq\int\limits\limits_{\mathcal{D}_{\Gamma_{0}}}\phi_{1}(\tau,z)\overline{\phi_{2}(\tau,z)}\,\eta^{k}\,e^{-\frac{4\pi my^{2}}{\eta}}\,\frac{\mathrm{d}\xi\wedge\mathrm{d}\eta\wedge\mathrm{d}x\wedge\mathrm{d}y}{\eta^{3}}\qquad(\phi_{1},\phi_{2}\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})).$
For $\phi\in J_{k,m}^{\mathrm{cusp}}(\Gamma_{0})$, one has the decomposition
(13) $\phi(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu}(\tau)\vartheta_{\mu,m}(\tau,z),$
where the function $h_{\mu}$ by Remark 2 on p. 287 of [Kr86] is a cusp form of weight $(k-\frac{1}{2})$ for the finite index subgroup
$\Gamma_{1}\coloneqq\Gamma_{0,1}(4m)=\bigg{\{}\begin{pmatrix}a&b\\ c&d\end{pmatrix}\in\Gamma_{0}\,\bigg{|}\,c\equiv 0,\,d\equiv 1\mod 4m\bigg{\}}\subseteq\Gamma_{0}(4m),$
and $\vartheta_{\mu,m}$ is the Jacobi theta function
@@
In fact, it is shown in Theorem 5.1 of [EZ85] that the decomposition (14) gives rise to the isomorphism
@@
$\phi_{1}(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu,1}(\tau)\vartheta_{\mu,m}(\tau,z)\quad\text{and}\quad\phi_{2}(\tau,z)=\sum_{\mu=0}^{2m-1}h_{\mu,2}(\tau)\vartheta_{\mu,m}(\tau,z)$